OpenStudy (anonymous):

Find the value of r so that the slope between the points is 4. (6, -2) and (r, -6) a. -5 b. 5 c. -42 d. 4 The slope of the line passing through (5, -4) and (5, -10) is ___ a. positive b. negative c. zero d. undefined If a 3 month old baby is 20 inches tall, and then at 7 months old, the baby is 26 inches tall, what is the average growth of the baby each month? a. 1.5 in. per month b. 2/3 in. per month c. 6 in. per month d. 6/7 in. per month

OpenStudy (onepieceftw):

1. the slope formula is\[\frac{ y_1-y_2 }{ x_1-x_2 }\] so plug in the known numbers. \[\frac{ -2-(-6) }{ 6-r }\] 6 minus what equals 1? Since the numerator is 4 to make the slope 4 the denominator needs to be 1.

OpenStudy (onepieceftw):

2. just use the slope formula I just showed you and you can determine if the slop is positive, negative, zero, or undefined. 3. The month is the x value and the age is the y value. So use the formula. \[\frac{ 20-26 }{ 3-7 }\] It'll be a fraction so just punch it into your calculator and you'll get the decimal equivalent.

OpenStudy (onepieceftw):

Hey can I get a medal? Or was I not helpful?